# Diffusion models for multivariate subsurface generation and efficient probabilistic inversion

### v3

@@ 2.4 Performance assessment
Prior modeling performance is assessed by comparing the unconditional subsurface realizations against the TI and realizations generated with two benchmark models, a GAN and a VAE, proposed by Miele et al. (2024) for an analogous case study. The statistical distance between the target and modeled distributions (either prior or posterior) were quantified using the root-mean-square-error (RMSE) and the Kullback-Leibler divergence (KLKL) (Kullback and Leibler, 1951). The similarity of spatial features in the realizations was assessed using the structural similarity index (SSIM) (Wang et al., 2004) defined as
@@ 3.1 Unconditonal modeling
The joint distributions of facies and co-located $I_{P}$ obtained from the sampled realizations (Fig. 2b, d, and f) show significantly superior modeling ability of the DM (Fig. 2b). The facies classes are modeled with negligible errors (i.e., facies values different from 0 and 1); the corresponding $I_{P}$ marginal distributions (plotted along the y-axes) match accurately the prior means and standard deviations, fitting the data with a KLKL divergence approximately five times lower than for GAN and VAE (Table 1). Further analysis of the networks unconditional modeling performances focusing on geostatistical metrics, are provided in Appendix A.
@@ 3.2.1 Linear conditioning results
For the linear conditioning case, we considered three test cases involving the contamination of $\bf{d_{obs}}$ (Figure 3b) with different levels of uncorrelated Gaussian noise ($\sigma_{Fac}$ and $\sigma_{I_{P}}$, see Table 2). Figure 5a shows the score magnitudes, expressed as L2-norms, for CDPS and DPS as a function of denoising/generative steps. For all settings, the scores of the data distribution have very similar values and decreasing trends; an average for the three noise levels is plotted as dashed lines. The likelihood scores show generally an inverse correlation between their magnitude and the data noise level, highlighting a stronger constraining effect of low-noise $\bf{d_{obs}}$. This correlation is negligible for the CDPS after 300 diffusion steps, with $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{d}|\textbf{x}_{t})$ being approximately 30 times lower than $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{x}_{t})$. A similar trend is observed in DPS only for the the largest data noise case (Data noise 3), while the likelihood score values generally increase with decreasing data error.
For each scenario, we generated 100 conditional samples. The convergence of the CDPS process is shown in Figure 5d as the WRMSE between the predicted $\hat{\textbf{x}}_{0}^{(t)}$ ($\hat{\bf{x}}_{0}^{(t)}=\textbf{x}_{0}$ at $t=0$) and $\bf{d_{obs}}$. Between 93 and 95% of the generated CDPS samples converged to a WRMSE <1.1<$1.$ For DPS, none of the samples converged when considering the lowest noise levels, and 99% of the samples converged to a WRMSE <1.1 for the two larger-error cases. Overall, the WRMSE of the DPS samples approach lower values faster than CDPS during the denoising process; for Data noise 1, the DPS samples ($\hat{\textbf{x}}_{0}^{(t)}$) initially overfit the data noise (WRMSE < 1) before diverging (WRMSE $\approx 1.5$). The WRMSE values obtained by CDPS also show comparatively larger variance (i.e., DPS samples follow similar paths).
@@ 3.2.2 Nonlinear inversion results
We ran the nonlinear inversion example for three noise-levels in $\bf{d_{obs}}$ (Fig. 3c), considering absolute and relative components of the standard deviations ($\sigma_{c}$ and $r$, respectively, in Table 2). Similar to the linear case, the prior scores (Fig. 5b) display a decreasing trend that does not change significantly for different noise levels. The $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{d}|\textbf{x}_{t})$ values of CDPS show a similar decreasing trend for the three noise levels, ranging from one-tenth to similar values as $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{x}_{t})$. Unlike the linear case, these scores show persistent differences between noise scales, with a maximum difference of one order of magnitude between the minimum and maximum data noise used. For DPS, the $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{d}|\textbf{x}_{t})$ values are comparatively larger than CDPS and rapidly exceed the prior scores, becoming up to ten times larger (Data noise 1).
@@
The CDPS has superior prediction accuracy. For the whole study area, the generated ensemble captures the true $I_{P}$ much better than DPS, as shown by the lower values of logS in Table 4 and in the trace example in Figure 8h (x = 25 m). Moreover, each realization reproduces the target $I_{P}$-distribution with a low KLKL divergence (Table 4) and presents fine-scale spatial variability patterns, which appear smoothed in DPS realizations (Fig. 8g).

### v4

@@ 2.4 Performance assessment
Prior modeling performance is assessed by comparing the unconditional subsurface realizations against the TI and realizations generated with two benchmark models, a GAN and a VAE, proposed by Miele et al. (2024) for an analogous case study. The statistical distance between the target and modeled distributions (either prior or posterior) were quantified using the root-mean-square-error (RMSE) and the Kullback-Leibler divergence ($KL$) (Kullback and Leibler, 1951). The similarity of spatial features in the realizations was assessed using the structural similarity index (SSIM) (Wang et al., 2004) defined as
@@ 3.1 Unconditonal modeling
The joint distributions of facies and co-located $I_{P}$ obtained from the sampled realizations (Fig. 2b, d, and f) show significantly superior modeling ability of the DM (Fig. 2b). The facies classes are modeled with negligible errors (i.e., facies values different from 0 and 1); the corresponding $I_{P}$ marginal distributions (plotted along the y-axes) match accurately the prior means and standard deviations, fitting the data with a $KL$ divergence approximately five times lower than for GAN and VAE (Table 1). Further analysis of the networks unconditional modeling performances focusing on geostatistical metrics, are provided in Appendix A.
@@ 3.2.1 Linear conditioning results
For the linear conditioning case, we considered three test cases involving the contamination of $\bf{d_{obs}}$ (Figure 3b) with different levels of uncorrelated Gaussian noise ($\sigma_{Fac}$ and $\sigma_{I_{P}}$, see Table 2). Figure 5a shows the score magnitudes, expressed as L2-norms, for CDPS and DPS as a function of denoising/generative steps. For all settings, the scores of the data distribution have very similar values and increasing trends; an average for the three noise levels is plotted as dashed lines. The likelihood scores show generally an inverse correlation between their magnitude and the data noise level, highlighting a stronger constraining effect of low-noise $\bf{d_{obs}}$. This dependence is negligible for the first half of the denoising process for CDPS, with $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{d}|\textbf{x}_{t})$ being around 30 times lower than $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{x}_{t})$. A similar ratio between prior and likelihood scores is observed in DPS for the the largest data noise case (Data noise 3), while the overall likelihood score values generally increase with decreasing data error. At the end of the denoising process, the likelihood scores are very similar for both CDPS and DPS. This is expected as the two formulations become identical when the denoiser error becomes negligible.
For each scenario, we generated 100 conditional samples. The convergence of the CDPS process is shown in Figure 5c as the WRMSE between the predicted $\hat{\textbf{x}}_{0}^{(t)}$ ($\hat{\bf{x}}_{0}^{(t)}=\textbf{x}_{0}$ at $t=0$) and $\bf{d_{obs}}$. Between 93 and 95% of the generated CDPS samples converged to a WRMSE <1.1<$1.$ For DPS, none of the samples converged when considering the lowest noise levels, and 99% of the samples converged to a WRMSE <1.1 for the two larger-error cases. Overall, the WRMSE of the DPS samples approach lower values faster than CDPS during the denoising process; for Data noise 1, the DPS samples ($\hat{\textbf{x}}_{0}^{(t)}$) initially overfit the data noise (WRMSE < 1) before diverging (WRMSE $\approx 1.5$). The WRMSE values obtained by CDPS also show comparatively larger variance (i.e., DPS samples follow similar paths).
@@ 3.2.2 Nonlinear inversion results
We ran the nonlinear inversion example for three noise-levels in $\bf{d_{obs}}$ (Fig. 3c), considering absolute and relative components of the standard deviations ($\sigma_{c}$ and $r$, respectively, in Table 2). Similar to the linear case, the prior scores (Fig. 5b) display an increasing trend that does not change significantly for different noise levels. The $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{d}|\textbf{x}_{t})$ values of CDPS show a rather similar increasing trend for the three noise levels, ranging from one-tenth to similar values as $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{x}_{t})$. Unlike the linear case, these scores show persistent differences between noise scales, with a maximum difference of one order of magnitude between the minimum and maximum data noise used. For DPS, the $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{d}|\textbf{x}_{t})$ values are comparatively larger than CDPS and rapidly exceed the prior scores, becoming up to ten times larger (Data noise 1).
@@
The CDPS has superior prediction accuracy. For the whole study area, the generated ensemble captures the true $I_{P}$ much better than DPS, as shown by the lower values of logS in Table 4 and in the trace example in Figure 8h (x = 25 m). Moreover, each realization reproduces the target $I_{P}$-distribution with a low $KL$ divergence (Table 4) and presents fine-scale spatial variability patterns, which appear smoothed in DPS realizations (Fig. 8g).